OpenStudy (anonymous):

\(\sf Simplify ~~\huge x^2-7x-\frac{8}{x^2}-4x-32\)

Nnesha (nnesha):

\[\huge\rm \frac{ x^2 -7x-8 }{ x^2 -4x-32 }\] factor both numerator and denominator

Nnesha (nnesha):

do you know how to factor quadratic equations ??

OpenStudy (anonymous):

FACTOR IT

OpenStudy (alexandervonhumboldt2):

nope

Nnesha (nnesha):

for first one ?? then nope

OpenStudy (alexandervonhumboldt2):

you need to use quadatic formula to find root so you can factor

Nnesha (nnesha):

okay lets factor first one \[\huge\rm x^2 -7x -8\]let me know what you get :)

Nnesha (nnesha):

find two number if you multiply them you should get -8 and if you add or subtract them you should get -7 :)

OpenStudy (anonymous):

-8 and 1

Nnesha (nnesha):

yes that is right :) :)

Nnesha (nnesha):

\[\huge\rm(x-8)(x+1)\] because leading coefficient is one so you can write like this ;)

Nnesha (nnesha):

\[\huge\rm x^2-4x -32\] what about this one ??

OpenStudy (anonymous):

-8 and 4?

Nnesha (nnesha):

yes right -8 times 4 = -32 and -8 +4 = -4

OpenStudy (anonymous):

Ah, thank you so much! Would you mind helping me with one more problem? :)

Nnesha (nnesha):

\[\frac{ (x-8)(x+1) }{ (x-8)(x+4) }\]

Nnesha (nnesha):

got it!! ??

Nnesha (nnesha):

my pleasure and SURE!!

OpenStudy (anonymous):

Yes, you're a great help!! And the other problem says to multiply, x^2/x+7 * x^2+11x+28/x^2+4x

Nnesha (nnesha):

\[ \huge\rm \frac{ x^2 }{ x+7 } \times \frac{ x^2 +11x +28 }{ x^2 +4x }\] O_O this is your question right ?

OpenStudy (anonymous):

Correct!

Nnesha (nnesha):

again factor both den. and numerator on right side

OpenStudy (anonymous):

Okay, how do you factor if there is only 2 numbers?

Nnesha (nnesha):

\[\frac{ x^2 +11x +28 }{ x^2 +4x }\] do you mean den. ? x^2 +4x ? this one ??

OpenStudy (anonymous):

Yes .

Nnesha (nnesha):

oaky so what is common in this both term x^2 and 4x ??

OpenStudy (anonymous):

I believe for the top it is 7 and 4? And the x.

Nnesha (nnesha):

yes right 7 times 4 = 28 7 +4 = 11 so yes

Nnesha (nnesha):

yes for the bottom part just x so what you left

OpenStudy (anonymous):

So would it be (x+7)(x+4) / (x+4)?

Nnesha (nnesha):

yes right but don't forget x which suppose to be at the bottom

OpenStudy (anonymous):

Okay, so then the two x+4 would cancel out.

Nnesha (nnesha):

\[\frac{ (x+7) (x+4)}{ x(x+4) }\] like this

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

Oh I see, so I canceled them out now I'm left with x^2/x+7 * x+7/x

Nnesha (nnesha):

yes right

OpenStudy (anonymous):

Okay, can you please help me with wht the next step is(:

Nnesha (nnesha):

what's the final answer ??

Nnesha (nnesha):

ohh okay i thought you said with next question

OpenStudy (anonymous):

Oh haha, I'm sorry!!

Nnesha (nnesha):

 you can cancel out x+7 ;)